Andretti Company has a single product called a Dak. The company normally produces and sells 60,000 Daks each year at a selling price of $32 per unit. The company’s unit costs at this level of activity are given below:

Direct materials. $10.00

Direct labor. 4.50

Variable manufacturing overhead. 2.30

Fixed manufacturing overhead. 5.00 ($300,000 total)

Variable selling expenses. 1.20

Fixed selling expenses. 3.50 ($210,000 total)

Total cost per unit. $26.50

Due to a strike in its supplier’s plant, Andretti Company is unable to purchase more material for the production of Daks. The strike is expected to last for two months. Andretti Company has enough material on hand to operate at 30% of normal levels for the two-month period. As an alternative, Andretti could close its plant down entirely for the two months. If the plant were closed, fixed manufacturing overhead costs would continue at 60% of their normal level during the two-month period and the fixed selling expenses would be reduced by 20% during the two-month period. Utilizing Excel, create a spreadsheet similar to Exhibit 12-1 on page 566 of the textbook to compare alternatives. Display your results using the total cost approach. Utilize formulas for all calculations and upload as an Excel document. Analyze the data you gathered to support your decision of whether the plant should continue operating or close for the two-month period. As part of your analysis please discuss other factors that could possibly influence your decision.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Andretti
Variable cost per unit Note
Direct Materials                10.00 A
Direct Labor                   4.50 B
Variable Manufacturing overhead                   2.30 C
Variable Selling Expenses                   1.20 D
Total Variable cost per unit                18.00 E=A+B+C+D
Sell Price Per unit                32.00 F
Contribution Per unit                14.00 G=F-E
Number of Units         60,000.00 H
Contribution amount      840,000.00 I=G*H
Fixed cost
Fixed manufacturing overhead      300,000.00 J
Fixed selling expenses      210,000.00 K
Plant close Amount $
Fixed manufacturing overhead at 60%         30,000.00 L=J/12*2*60%
Fixed selling expenses at 80%         28,000.00 M=K/12*2*80%
Total cost incurred if plant closed        58,000.00
30% capacity Amount $
Number of units           3,000.00 N=H/12*2*30%
Contribution Per unit                14.00 See G
Contribution earned        42,000.00 O=N*G
Less: Fixed manufacturing overhead         50,000.00 P=J/12*2
Fixed selling expenses         35,000.00 Q=K/12*2
Net Income      (43,000.00) R=O-P-Q
Decrease in loss by        15,000.00 S=M+R
So Andretti should not close the plant for two months but operate it at 30% capacity.
Other factor that could possibly influence the decision are:
The assumptions may not be correct. Company may incur costs then estimated. It will impact the decision.